OpenStudy (anonymous):

Please help. Use basic identities to simplify the expression. cosecant of theta times cotangent of theta divided by secant of theta.

OpenStudy (lynfran):

\[\frac{ \csc \theta*\cot \theta }{ \sec \theta }?\]

OpenStudy (anonymous):

Yes

OpenStudy (lynfran):

well all you need to do here is \[\sec \theta=\frac{ 1 }{ \cos\theta}\]then multiple that to simplify

OpenStudy (lynfran):

so we get \[\frac{ \csc \theta*\cot }{ \frac{ 1 }{ \cos \theta } }\]

OpenStudy (anonymous):

sec2θ cot2θ csc2θ 1

OpenStudy (anonymous):

Is it the second one?